OpenStudy (anonymous):

Find the properties for the ellipse with the equation 36x^2 + 4y^2 = 9.

OpenStudy (anonymous):

I am asked to find the domain. Im sure its simple i just need to know what the next step is.

OpenStudy (anonymous):

@LastDayWork

OpenStudy (lastdaywork):

Domain implies the acceptable values of x and y. Use the fact that the square of a real number cannot be negative; can you tell me the maximum and minimum possible value for x (forget y for the time being)

OpenStudy (anonymous):

Im not sure.. D:

OpenStudy (anonymous):

I thought that you weer supposed to divide the whole equation by 9 to make teh right side of the equation equal 1.

OpenStudy (lastdaywork):

To find the maximum possible value of x^2 ; substitute y^2 with the minimum possible value. What do you think is the minimum value for y^2 ??

OpenStudy (anonymous):

Im guessing one. How would you find it?

OpenStudy (lastdaywork):

Minimum value for y^2 = 0 (as it cannot be negative). Now put y^2 = 0 in the original equation and solve for x

OpenStudy (lastdaywork):

BTW do you know how to solve inequalities ??

OpenStudy (anonymous):

x = 1/2: and give an example of an inequality (im not so good on names)

OpenStudy (anonymous):

i am also supposed to find E so i would need to find c and a

OpenStudy (lastdaywork):

Can you solve for x \[36x^2 \le 9\]

OpenStudy (anonymous):

not sure. i get x= plusorminus 1/2 at first glance but...

OpenStudy (lastdaywork):

So you never solved such equations, right ??

OpenStudy (anonymous):

i have. I just haven't in a while

OpenStudy (lastdaywork):

The answer is \[x \epsilon (-1/2 , 1/2)\] You just reported it in the wrong way. :)

OpenStudy (lastdaywork):

Now, similarly find the domain of y.

OpenStudy (anonymous):

haha alright. I just turned in the assignment with some other questions about that problem and got a 100% thanks for the help

OpenStudy (lastdaywork):

Actually a correction; \[x \epsilon \left[ \frac{ -1 }{ 2 },\frac{ 1 }{ 2 } \right]\]

OpenStudy (lastdaywork):

:)